Discuss with Askiitians Tutors

if 1/1^4+1/2^4+1/3^4......=pie^4/90
then find answer for 1/1^4+1/3^4+1/5^4 .....

gunjan
11 Years agoGrade

1 Answer

Nishant Vora
11 Years ago
Hello student, Please find the solution

[1/1^4+1/2^4+1/3^4 +1/4^4+1/5^4+1/6^4......=\pi ^4/90]

[(1/1^4+1/3^4+1/5^4 + .....)+(1/2^4+1/4^4+1/6^4+......)=\pi ^4/90]

[(1/1^4+1/3^4+1/5^4 + .....)+1/2^4 (1/1^4+1/2^4+1/3^4+......)=\pi ^4/90]

[(1+1/2^4)(1/1^4+1/2^4+1/3^4+......)=\pi ^4/90]
[(1/1^4+1/2^4+1/3^4+......)=\pi ^4/90*\frac{1}{(1+1/2^4)}]

Now you can simplyfy